OpenStudy (anonymous):

Solve. Write the answer in simplest form. (Must show work!) 1. d - 2/5 = 3/10

OpenStudy (mertsj):

1. Add 2/5 to both sides.

OpenStudy (mathlegend):

Can you do that @connor_ ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I just need the answer and the work so I can understand.

OpenStudy (anonymous):

d-(2/5)=(3/10) d-(4/10)=(3/10) add 4/10 to both sides d=(4/10 + 3/10) d=(7/10)

OpenStudy (anonymous):

Thank u so much! i understand now!